# DitHub: A Modular Framework for Incremental Open-Vocabulary Object Detection

### v1

@@ 5 Experimental Studies
This section evaluates DitHub in two parts. The first assesses its adaptability to multiple tasks using the Incremental Vision-Language Object Detection (IVLOD) protocol [2] in Sec. 5.1. The second, in Sec. 5.2, examines the method’s specialization capabilities. The section ends with ablation studies in Sec. 5.3.
@@ 5.1 Incremental Vision-Language Object Detection
Table 1: Comparison of mAP values across ODinW-13 tasks. Avg represents the average across the $13$ tasks, while zero-shot performance is reported in the ZCOCO column. Best results are highlighted in bold.
Evaluation settings. Following [2], we employ ODinW-13 [17] to assess the model ability to learn new tasks while mitigating forgetting of previously seen data. ODinW-13 consists of $13$ sub-datasets, spanning from the traditional Pascal VOC [6] to more challenging domains with significant distribution shifts, such as thermal imagery. Also, the IVLOD setting mandates that the Open-Vocabulary detector maintains robust zero-shot capabilities after fine-tuning. These capabilities are tested on MS COCO [19].
One issue of ODinW-13 is that most classes are confined to a single task among the $13$. However, DitHub is conceived to handle classes that reoccur across different domains, necessitating updates to their respective modules. To provide an additional perspective on incremental object detection and model reuse, we introduce a variant of ODinW-13, referred to as ODinW-O. Starting from ODinW-35, we construct ODinW-O by retaining only the classes that are shared across tasks while discarding those that appear in only one task. Further details are provided in App. B.
For comparisons, we employ mean Average Precision (mAP) and compare DitHub against six state-of-the-art methods. CL-DETR [22] and OW-DETR [9] are explicitly designed for incremental learning, while TFA [43] specializes in few-shot adaptation. AT [10] has been extended to both incremental and few-shot learning, whereas iDETR [4] integrates both aspects. Finally, ZiRa [2], our primary competitor, is specifically tailored for IVLOD.
Implementation details. We focus on Grounding DINO and adapt only its encoder network, as further validated in Sec. A.1. The merging coefficient $\lambda$ for the $A$ matrix is set to $0.3$ for ODinW-13 and $0.1$ for ODinW-O; for the matrix $B$, we set $\lambda=0.7$. The choice of $\lambda$ and its underlying rationale are further discussed in Sec. 5.3. During training, we allocate an equal number of epochs to the warmup and the subsequent specialization phases — App. C for further hyperparameter details.
Performance on ODinW-13. Table 1 presents the IVLOD results, including the average across the $13$ tasks (Avg) and zero-shot preservation (ZCOCO). We evaluate performance in both the standard setting with the full dataset and the few-shot scenarios, considering 1, 5, and 10 shots per class. All baseline results, except those of Grounding DINO, ZiRa, and our method, are taken directly from [2]. Each number represents the average of three runs with different random seeds, with standard deviations provided in Sec. D.1.
In the full-shot setting, DitHub outperforms its main competitor, ZiRa, by a substantial + 4.21 mAP points and achieves the best results in nine out of thirteen tasks. Notably, on ZCOCO, our approach outperforms ZiRa by a mAP gain of + 0.75, setting a new state-of-the-art in both incremental and zero-shot retention capabilities. Although DitHub is not explicitly designed for data-scarce conditions, DitHub achieves state-of-the-art performance even in the few-shot setting. In particular, in the five-shot and ten-shot settings, DitHub emerges as the best-performing approach, with the ten-shot scenario showing a slightly larger improvement over ZiRa (+ 1.23 mAP). Other baselines struggle to remain competitive in this regime. Our method also performs well in the one-shot setting, where iDETR achieves the highest average performance, followed closely by DitHub and ZiRa.
These findings — optimal performance in incremental learning, effective retention of zero-shot capabilities, and robustness in the low-data regime — clearly highlight the advantages of our modular approach, with the advanced update system that will be dissected in the analyses below.
ZCOCO Avg Ae Hw Pv Sd Th Ve
@@
ZiRa 44.43 57.63 39.92 68.00 64.90 46.26 77.26 49.47
DitHub 46.51 62.38 53.35 71.07 71.01 41.75 80.21 56.90
Table 2: mAP values on ODinW-O. Best in bold.
@@
Table 3: Experts not Experts (EneE). Training by selecting a random expert.
Performance on ODinW-O. The ODinW-O scenario emphasizes that frequently encountered classes may reappear in subsequent tasks or at later times, such as the class “person” in common industrial and smart-city scenarios. As detailed in Sec. B.2, ODinW-O guarantees that each class is revisited at least once in subsequent tasks. In this benchmark, we primarily compare our approach with ZiRa, as it is the most competitive method in the standard IVLOD setting, excluding ours. The results in Tab. 2 indicate that our approach secures a substantial improvement of + 4.75 in mAP and a + 2.08 mAP gain on ZCOCO over ZiRa. We attribute the success of our method to its class-oriented modular design, which enables selective updates to recurring concepts. This design effectively prevents the model from inadvertently overwriting crucial knowledge associated with other classes, tasks, or domains.
5.2 Evaluation of Specialized Detection Modules
Here, we delve into the properties of our specialization framework and the benefits of composing class-specific experts. To this aim, we introduce an ablative approach called Experts-not-Experts (EnE), which emphasizes the absence of class specialization in our experts. EnE is identical to DitHub, except during training, where it does not employ the stochastic strategy introduced in Sec. 4.1 and designed to specialize the experts. Instead, given an image from the current task, the EnE approach performs the training step on one module chosen at random among the classes of the entire task. As a result, modules no longer specialize in a class but instead learn from the whole task data, including images without their designated class. Consequently, they become general task learners rather than class-specific experts.
Figure 4: Specialization capabilities for rare classes of ODinW-13.
In Tab. 3, we offer a comparative analysis between our DitHub and its non-specialized variant, i.e., EnE. Notice that, at inference time, EnE acts as an ensemble of modules, each trained on the same data but with a different random ordering. In contrast, DitHub employs class-specialized modules with an expertise on data containing the same class. This targeted specialization improves performance, as shown in Tab. 3, where per-class specialization results in a + 1.23 mAP improvement on average. Additionally, albeit to a lesser extent, the ZCOCO metric still benefits from an improvement. These results affirm that a system composed of modular experts, each a custodian of distinct knowledge, offers a promising approach for systems requiring continuous and selective adaptation.
@@ 5.2 Evaluation of Specialized Detection Modules
Evaluation on rare classes. As discussed in Section 4.3, DitHub facilitates fine-grained control during inference, enabling users to tailor the system to their specific use case and deploy the most appropriate modules. This includes the option to use a single module specialized on a target class. To assess single-module effectiveness, we conduct a targeted evaluation on rare classes from ODinW-13, activating only the corresponding target class module. Results in Fig. 4 show that DitHub consistently achieves the best results across rare classes, highlighting the benefits of class specialization. EnE and DitHub outperform the zero-shot Grounding DINO baseline, though EnE underperforms for the Starfish category.
Figure 5: Showing the natural unlearning capabilities of DitHub.
@@
Specialized modules allows training-free unlearning. The advantages of DitHub extend beyond additive incremental knowledge compositionality. Indeed, we can show that expert modules can be useful to undermine or remove some specific knowledge or content from a model, by simply subtracting the corresponding weights from the base model. This operation, which resembles the concept of zero-training unlearning, has been explored by the authors of [13, 28] in the context of image and text categorization tasks; to the best of our knowledge, we are the first to explore its application within the context of object detection.
@@
Specifically, we test whether class knowledge can be unlearned from the zero-shot Grounding DINO baseline by subtracting the corresponding class module from the pre-trained model: $\mathbf{W}^{\prime}=\mathbf{W}-\alpha BA_{c}$, where we set $\alpha=0.3$. As a result, we provide a qualitative visualization shown in Fig. 5: on the left, the zero-shot model correctly recognizes common classes, while on the right, applying the negative module of a specific class reduces its mAP, effectively unlearning it and leading to false negatives. Additionally, we provide a quantitative evaluation of the unlearning process in Sec. D.2.
5.3 Ablation Studies
Figure 6: Sensitivity analysis on $\lambda_{A}$ and $\lambda_{B}$. Performance measured as harmonic mean of Avg mAP (across tasks) and ZCOCO.
@@ 5.3 Ablation Studies
Sensitivity Analysis of $\lambda.$ Analyzing the variation of the merging coefficients $\lambda_{A}$ and $\lambda_{B}$ offers valuable insights into our methodology. To show that, we define an aggregate metric that is the harmonic mean between DitHub’s full-shot performance on ZCOCO and Avg, ensuring a balanced evaluation of specialization (Avg) and zero-shot generalization (ZCOCO). Figure 6 presents the performance trends for ODinW-13 (top) and ODinW-O (bottom) while varying $\lambda_{A}$ and $\lambda_{B}$ in increments of $0.2$.
@@
Considering $\lambda_{A}$, its role is significant while merging the expert A matrices (see Eq. 3). From Fig. 6, lower values of $\lambda_{A}$ lead to better results, indicating that prioritizing past expert knowledge over the warmup module is more effective. Interestingly, while $\lambda_{A}=0.3$ performs best for ODinW-13, ODinW-O exhibits a monotonic decline as $\lambda_{A}$ increases. This contrast arises from dataset structures: ODinW-O consists solely of overlapping classes across tasks, where retaining past expertise is critical, whereas ODinW-13 features diverse class distributions per task, requiring greater integration of domain-specific warmup knowledge.
@@
On Warmup and Shared B. To evaluate the impact of each component in DitHub, we progressively integrate them and present the results in Fig. 7. We start with the zero-shot performance of Grounding DINO (GD) as a reference, followed by a baseline fine-tuning procedure (Base in Fig. 7), where an $A$ matrix is trained for each class without merging strategies, omitting the warmup phase and using a shared $B$ matrix. This method leads to complete forgetting as new tasks are introduced, but still outperforms Grounding DINO, which struggles with pretraining knowledge alone.
Building on this baseline, we first incorporate Eq. 4, but see no significant performance gain. This is expected, as the shared $B$ matrix only represents a tiny portion of the total parameters. Without class-specific focus, Eq. 4 alone does not enhance retention. However, adding the warmup phase (+wu) provides a clear improvement, as it stabilizes training and helps rare classes by offering a robust initialization.
When we further include Eq. 4 in the warmup-enhanced fine-tuning strategy, a modest improvement is observed. However, the most substantial performance gain occurs when Eq. 3 is also integrated into the methodology (DitHub in Fig. 7). This highlights the importance of preserving past domain knowledge within class-specific modules, ensuring a more stable and effective incremental learning process. Notably, the combination of all three key components — the warmup phase for reliable initialization, the merging strategy for $A$ matrices to retain class-specific knowledge across domains, and the shared $B$ matrix approach for efficiency — results in a significant + 6.22 improvement over the baseline where none of these strategies are applied.
@@ 6 Conclusions
Looking ahead, we aim to further leverage DitHub’s modular library management to deepen our understanding of knowledge transfer between efficient modules for object detection. For instance, if our library contains an “axolotl” module trained on real-world images and we possess domain knowledge for the cartoon image domain, can we synthesize a module specialized in detecting cartoon “axolotls” through knowledge transfer? DitHub will provide a fork of this manuscript to explore such questions.
Supplementary Material Overview
This supplementary material complements the main manuscript by providing additional analyses, insights, and results on DitHub. Specifically, we delve deeper into efficiency considerations of our adaptation approach (App. A), highlighting blockwise adaptation and LoRA specializations. We introduce the novel ODinW-O benchmark for incremental learning with overlapping classes and domain shifts (App. B), and present insights on unlearning class-specific knowledge (Sec. D.2). Additionally, we clarify critical implementation specifics (App. C), present extensive robustness checks (Sec. D.1), and showcase qualitative examples (App. E) to illustrate the practical benefits of our methodology.
Appendix A Further Insights on Efficiency
A.1 Blockwise Adaptation Analysis
While Grounding DINO consists of four main components — an image backbone for visual encoding, a textual backbone for language understanding, and an encoder-decoder pair for detection — we focus our adaptation on the latter two. This choice is motivated by the need for a more detection-centric adaptation and efficiency considerations, as the textual backbone, being a BERT Base model [3], accounts for more than half of Grounding DINO ’s parameters. In Table B, we present results for adapting only the encoder (Figure A), only the decoder (Figure A), and both encoder and decoder (Figure A). DitHub employs the first approach — adapting only the encoder — as it delivers the best performance on both ZCOCO and Avg metrics.
A.2 Employing Full LoRA for Class-Specificity
While our methodology leverages a shared and unique $B$ matrix for all tasks, Tab. A presents results obtained using class-specific pairs. In this variant of DitHub, both the $A$ and $B$ low-rank matrices of LoRA undergo the same specialized optimization process: an initial warmup phase followed by dedicated specialization. In other words, the training pipeline for matrix $B$ mirrors that of matrix $A$ described in Section 4.1. Applying our full set of strategies to both matrices yields a + 1.03 improvement in average performance across the tasks of ODinW-13, albeit at the cost of doubling the number of trainable parameters (Fig. 3). Prioritizing efficiency, we opt for a lightweight implementation that trades a marginal performance gain for significantly reduced memory requirements. Notably, ZCOCO results favor our efficient approach, as a single, unspecialized $B$ matrix remains more aligned with the pretraining distribution, thereby mitigating degradation in zero-shot capabilities.
Method ZCOCO Avg
DitHub 47.01 62.19
DitHub (LoRA full) 46.66 63.22
Table A: LoRA full results on ODinW-13, with DitHub results taken from Table 1 for reference.
(a) Encoder
(b) Decoder
(c) Decoder/Encoder
Figure A: Blockwise analysis of adaptation strategies: (a) Adaptation of the encoder only (our chosen approach), (b) Adaptation of the decoder only and (c) Joint adaptation of both the decoder and encoder.
@@ A.2 Employing Full LoRA for Class-Specificity
Table B: Performance comparison across datasets when adapting only the decoder (”Only Dec”), both encoder and decoder (”Dec/Enc”), and encoder only (DitHub).
Appendix B Additional Benchmark Details

### v2

@@ 5.1 Incremental Vision-Language Object Detection
Evaluation settings. Following [4], we employ ODinW-13 [20] to assess the model ability to learn new tasks while mitigating forgetting. ODinW-13 includes $13$ sub-datasets, spanning from the traditional Pascal VOC [8] to more challenging domains with significant distribution shifts, such as thermal imagery. Also, the IVLOD setting requires the Open-Vocabulary detector to retain strong zero-shot capabilities after fine-tuning, which are evaluated on MS COCO [22].
One issue of ODinW-13 is that most classes are confined to a single task among the $13$. However, DitHub is conceived to handle recurring classes across domains, necessitating updates to their respective modules. To better assess incremental detection and module reuse, we introduce ODinW-O (Overlapped), a subset of ODinW-35 that is designed to reflect realistic scenarios where classes can reoccur across tasks over time. Further details are provided in App. C.
For comparisons, we employ mean Average Precision (mAP) and compare DitHub against six state-of-the-art methods. CL-DETR [25] and OW-DETR [11] are explicitly designed for Incremental Learning, while TFA [48] specializes in few-shot adaptation. AT [12] has been extended to both incremental and few-shot learning, whereas iDETR [6] natively integrates both aspects. Finally, ZiRa [4], our primary competitor, is specifically tailored for IVLOD.
Implementation details. We focus on Grounding DINO and adapt only its encoder network, as further validated in Sec. A.1. The merging coefficient $\lambda$ for the $A$ matrix is set to $0.3$ for ODinW-13 and $0.1$ for ODinW-O; for the matrix $B$, we set $\lambda=0.7$. The choice of $\lambda$ and its underlying rationale are further discussed in Sec. 5.2. During training, we allocate an equal number of epochs to the warmup and the subsequent specialization phases — see App. E for further hyperparameter details.
Performance on ODinW-13. Tab. 1 presents the IVLOD results, including the average across the $13$ tasks (Avg) and zero-shot preservation (ZCOCO). We evaluate performance in both the full-data setting and few-shot scenarios with $1$, $5$, and $10$ shots per class. All baseline results, except those of Grounding DINO, ZiRa, and our method, are taken directly from [4]. Each number represents the average of three runs with different random seeds, with standard deviations provided in Sec. D.1.
Table 1: Comparison of mAP values across ODinW-13 tasks. Avg represents the average across the $13$ tasks, while zero-shot performance is reported in the ZCOCO column. Bold indicates best results.
@@
DitHub $\boldsymbol{47.01}$ $\boldsymbol{62.19}$ $\boldsymbol{34.62}$ $\boldsymbol{50.65}$ $70.46$ $68.56$ $49.28$ $\boldsymbol{65.57}$ $\boldsymbol{69.58}$ $71.10$ $\boldsymbol{56.65}$ $\boldsymbol{70.88}$ $\boldsymbol{52.82}$ $\boldsymbol{79.30}$ $\boldsymbol{68.18}$
In the full-shot setting, DitHub outperforms ZiRa by a substantial $\boldsymbol{\texttt{+}4.21}$ mAP and achieves the best results in nine out of thirteen tasks. Notably, on ZCOCO, it surpasses ZiRa by a $\boldsymbol{\texttt{+}0.75}$ mAP, setting a new state-of-the-art in both incremental and zero-shot retention. While not explicitly designed for data-scarce conditions, DitHub also excels in the few-shot setting. In particular, in the five-shot and ten-shot settings, DitHub emerges as the best-performing approach, with the ten-shot scenario showing a slightly larger improvement over ZiRa ($\boldsymbol{\texttt{+}1.23}$ mAP). Our method also performs well in the one-shot setting, where iDETR leads, followed closely by DitHub and ZiRa.
These findings – optimal performance in Incremental Learning, effective retention of zero-shot capabilities, and robustness in the low-data regime – clearly highlight the advantages of our modular approach.
Performance on ODinW-O. The ODinW-O scenario emphasizes that frequently encountered classes may reappear in subsequent tasks or at later times, such as the class “person” in common industrial and smart-city scenarios. As detailed in Sec. C.2, ODinW-O guarantees that each class is revisited at least once in subsequent tasks. In this benchmark, we primarily compare our approach with ZiRa, as it is the most competitive method in the standard IVLOD setting, excluding ours.
Table 2: mAP values on ODinW-O. Best in bold.
Method ZCOCO Avg Ae Hw Pv Sd Th Ve
@@
ZiRa $44.43$ $57.63$ $39.92$ $68.00$ $64.90$ $\boldsymbol{46.26}$ $77.26$ $49.47$
DitHub $\boldsymbol{46.51}$ $\boldsymbol{62.38}$ $\boldsymbol{53.35}$ $\boldsymbol{71.07}$ $\boldsymbol{71.01}$ $41.75$ $\boldsymbol{80.21}$ $\boldsymbol{56.90}$
The results in Tab. 2 indicate that our approach secures a substantial improvement of $\boldsymbol{\texttt{+}4.75}$ in mAP and a $\boldsymbol{\texttt{+}2.08}$ mAP gain on ZCOCO over ZiRa. We attribute the success of our method to its class-oriented modular design, which enables selective updates to recurring concepts. This design prevents unintentional overwriting of knowledge across classes, tasks, and domains.
Figure 4: Normalized Average mAP as tasks progress. To assess performance degradation, we normalize the mAP for each task by dividing the current value by the original mAP recorded at the end of that task, i.e., if the curve $\approx 0.90$, the model retains $90\%$ of its original performance.
Performance analysis as tasks progress. In Fig. 4, we present an incremental analysis comparing the performance of DitHub and its strongest competitor, ZiRa, after each task. The evaluation, conducted on both ODinW-13 (left) and ODinW-O (right), allows us to assess the retention of both zero-shot and in-domain capabilities throughout the task sequence, providing a direct measure of their resistance to forgetting (further discussed in App. B). On ZCOCO, DitHub exhibits stable performance across tasks, with a negligible loss in zero-shot capabilities — unlike ZiRa, which shows a more pronounced decline. When evaluating performance across the sequence of downstream tasks, both methods demonstrate a decreasing trend, indicating that forgetting remains a significant challenge in Incremental Object Detection. Nevertheless, as more tasks are introduced, the performance gap widens, with DitHub consistently outperforming its competitor. This highlights the effectiveness of our modular design, which enables scalable knowledge organization and adaptation as the number of tasks grows. Moreover, the gap is still present in the ODinW-O setting introduced in this work, which models the realistic scenario where subsequent tasks may share classes. This setting, being inherently more susceptible to forgetting, further underscores the advantages of our approach.
5.2 Model Analysis and Ablation Studies
We examine the properties of our specialization framework and the benefits of composing class-specific experts. To this aim, we introduce an ablative approach, Experts-not-Experts (EnE), which removes class specialization. EnE is identical to DitHub, except during training, where it omits the stochastic strategy from Sec. 4.1. Instead, given an image from the current task, EnE performs the training step on one module chosen at random among all task classes. As a result, modules no longer specialize in a class but instead learn from the whole task data, including images without their designated class. Consequently, they become generalist learners rather than class-specific experts.
In Tab. 3, we compare DitHub with its non-specialized variant, i.e., EnE. Notice that, at inference, EnE acts as an ensemble of class-agnostic modules. In contrast, DitHub employs class-specialized counterparts with an expertise on data containing the same class. This targeted specialization yields a $\boldsymbol{\texttt{+}1.23}$ mAP improvement on average, with additional gains also on the ZCOCO metric. These results affirm that a system of modular experts, each with distinct knowledge, offers a promising approach for systems requiring continuous and selective adaptation.
Table 3: Experts-not-Experts (EnE). Training by selecting a random expert.
@@ 5.2 Model Analysis and Ablation Studies
Figure 5: Results on rare classes of ODinW-13.
Figure 6: Showing unlearning skills of DitHub.
@@
Evaluation on rare classes. As discussed in Sec. 4.3, DitHub offers fine-grained control at inference, enabling users to tailor the system to their specific use case — including the option to use a single module specialized for a target class. To assess this, we conduct a targeted evaluation on rare classes from ODinW-13, activating only the corresponding target class module. Results in Fig. 6 show that DitHub achieves the best results on rare classes, highlighting the benefits of class specialization.
@@
Specialized modules allows training-free unlearning. DitHub’s advantages go beyond additive Incremental Learning. Expert modules can also be used to remove specific knowledge by simply subtracting the corresponding weights from the base model. This operation, akin to zero-training unlearning, has been studied in image and text categorization [15, 33]; to the best of our knowledge, we are the first to explore it within the context of Object Detection. For example, this enables the model to unlearn the detection of sensitive object categories – such as faces, license plates, or medical equipment – when required to comply with privacy regulations or data retention policies.
@@
Specifically, we test whether class knowledge can be unlearned from the zero-shot Grounding DINO baseline by subtracting the corresponding class $c$ module from the pre-trained model: $W^{\prime}=W-\alpha BA_{c}$, whit $\alpha=0.3$. Fig. 6 illustrates this qualitatively: on the left, the zero-shot model correctly detects common classes, while on the right, subtracting a class module reduces its mAP, effectively unlearning it and leading to false negatives. A quantitative analysis is provided in Sec. D.2.
@@
Sensitivity Analysis of $\boldsymbol{\lambda}$. Varying the merging coefficients $\lambda_{A}$ and $\lambda_{B}$ offers valuable insights into our method. To this end, we define an aggregate metric as the harmonic mean of DitHub’s full-shot performance, balancing specialization (Avg) and zero-shot generalization (ZCOCO). Fig. 8 shows performance trends for ODinW-13 (top) and ODinW-O (bottom) as $\lambda_{A}$ and $\lambda_{B}$ in $0.2$ increments.
@@
Considering $\lambda_{A}$, its role is significant while merging the expert $A$ matrices (see Eq. 2). From Fig. 8, lower values of $\lambda_{A}$ lead to better results, indicating that prioritizing past expert knowledge over the warmup module is more effective. Interestingly, while $\lambda_{A}=0.3$ performs best for ODinW-13, ODinW-O exhibits a monotonic decline as $\lambda_{A}$ increases. The contrast reflects dataset structure: ODinW-O includes overlapping classes, emphasizing knowledge retention, whereas ODinW-13 has varied class distributions, demanding stronger integration of domain-specific warmup knowledge.
@@
Figure 7: Sensitivity analysis on $\lambda_{A}$ and $\lambda_{B}$. Performance measured as harmonic mean of Avg and ZCOCO mAP.
@@
On Warmup and Shared $\boldsymbol{B}$. To evaluate the impact of DitHub’s components, we progressively integrate them and present results in Fig. 8. We establish the zero-shot performance of Grounding DINO (GD) as a reference. Our baseline fine-tuning procedure (Base) trains class-specific $A$ matrices and the shared $B$ matrix independently for each new task, omitting both parameter merging and the warmup phase. While this baseline suffers from complete catastrophic forgetting, it still demonstrates improved performance compared to the zero-shot Grounding DINO.
Building on this baseline, we first add the warmup phase (+wu), which yields a notable improvement by stabilizing training and facilitating the learning of rare classes through robust initialization. Subsequently, incorporating Eq. 3 into the warmup-augmented strategy results in a modest improvement — this is expected, as the shared $B$ matrix constitutes only a small fraction of the total parameters. Crucially, the most substantial performance gain is achieved upon integrating Eq. 2 (DitHub in Fig. 8), highlighting the importance of retaining domain-specific knowledge within class-specific modules to support stable and effective Incremental Learning. Collectively, the synergistic combination of these three key components – the warmup phase for robust initialization, the merging strategy for $A$ matrices to retain class-specific knowledge across domains, and the shared $B$ matrix for efficiency – yields a substantial $\boldsymbol{\texttt{+}6.13}$ mAP improvement compared to the baseline model.
@@ 6 Conclusions
In future works, we aim to further leverage modular library management to deepen our understanding of knowledge transfer between efficient modules for Object Detection. For instance, if our library contains an “axolotl” module trained on real-world images and we possess domain knowledge for the cartoon image domain, can we synthesize a module specialized in detecting cartoon “axolotls” through knowledge transfer? DitHub will provide a fork of this manuscript to explore such questions.
Appendix Overview
This Appendix provides supplementary information regarding the main manuscript.
  • App. A — Further Insights on Efficiency: Optimizes DitHub’s performance when adapting Grounding DINO.
    • Sec. A.1 — Blockwise Adaptation Analysis: Justifies adapting only the encoder for DitHub based on empirical metrics.
    • Sec. A.2 — Employing Full LoRA for Class-Specificity: Explores class-specific LoRA, favoring a parameter-efficient shared $B$ matrix.
  • App. B — Further Insights on Forgetting: Analyzes DitHub’s anti-forgetting capabilities.
    • Sec. B.1 — Forgetting on ODinW-13: Presents anti-forgetting results for ODinW-13.
    • Sec. B.2 — Forgetting on ODinW-O: Demonstrates robust anti-forgetting on the ODinW-O benchmark.
  • App. C — Additional Benchmark Details: Provides comprehensive dataset information.
    • Sec. C.1 — ODinW-13 and MS COCO: Describes ODinW-13 for specialization and MS COCO (ZCOCO) for zero-shot evaluation.
    • Sec. C.2 — ODinW-O: Introduces the ODinW-O benchmark for Incremental Learning and domain shifts.
  • App. D — Additional Results: Presents complementary experimental information.
    • Sec. D.1 — Standard Deviations: Reports standard deviations for reproduced methods from Tab. 1 and Tab. 2.
    • Sec. D.2 — Unlearning Additionals: Illustrates DitHub’s modular capability for selective knowledge removal (Fig. C).
  • App. E — Implementation Details: Outlines experimental setup, including hardware, model configurations, and optimization specifics.
  • App. F — Limitations and Broader Impact: Discusses current limitations and broader societal implications.
  • App. G — Qualitative Results: Provides visual comparisons (Fig. D) demonstrating DitHub’s improved detection and robustness.
Table A: Performance comparison across datasets when adapting only the decoder (Only Dec), both encoder and decoder (Dec/Enc), and encoder only (DitHub).
@@
(a) Encoder
(b) Decoder
(c) Decoder/Encoder
Figure A: Blockwise analysis of adaptation strategies: (a) encoder-only adaptation (our selected method), (b) decoder-only adaptation, and (c) joint encoder-decoder adaptation.
Appendix A Further Insights on Efficiency
A.1 Blockwise Adaptation Analysis
While Grounding DINO consists of four main components – an image backbone for visual encoding, a textual backbone for language understanding, and an encoder-decoder pair for detection – we focus our adaptation on the latter two. This choice is motivated by the need for a more detection-centric adaptation and efficiency considerations, as the textual backbone, being a BERT Base model [5], accounts for more than half of Grounding DINO ’s parameters. In Tab. A, we present results for adapting only the encoder (Fig. A), only the decoder (Fig. A), and both encoder and decoder (Fig. A). DitHub employs the first approach – adapting only the encoder – as it delivers the best performance on both ZCOCO and Avg metrics.
A.2 Employing Full LoRA for Class-Specificity
Table B: LoRA full results on ODinW-13, with DitHub results taken from Tab. 1 for reference.
Method ZCOCO Avg
DitHub $47.01$ $62.19$
DitHub (LoRA full) $46.66$ $63.22$
While our methodology leverages a shared and unique $B$ matrix for all tasks, Tab. B presents results obtained using class-specific pairs. In this variant of DitHub, both the $A$ and $B$ low-rank matrices of LoRA undergo the same specialized optimization process: an initial warmup phase followed by dedicated specialization. In other words, the training pipeline for matrix $B$ mirrors that of matrix $A$ described in Sec. 4.1. Applying our full set of strategies to both matrices yields a $\boldsymbol{\texttt{+}1.03}$ mAP improvement in average performance across the tasks of ODinW-13, albeit at the cost of doubling the number of trainable parameters (Fig. 3). Prioritizing efficiency, we opt for a lightweight implementation that trades a marginal performance gain for significantly reduced memory requirements. Notably, ZCOCO results favor our efficient approach, as a single, unspecialized $B$ matrix remains more aligned with the pretraining distribution, thereby mitigating degradation in zero-shot capabilities.
Appendix B Further Insights on Forgetting
This section presents an additional analysis of the anti-forgetting properties of DitHub, complementing the study in Sec. 5.2 by offering an alternative perspective. We report the forgetting metric for both ODinW-13 and ODinW-O, computed for DitHub and its primary competitor, ZiRa [4]. This metric is calculated using the canonical definition [3, 1]: the difference in performance on a specific task after learning the entire sequence of tasks compared to the original performance on that task immediately after its initial learning.
B.1 Forgetting on ODinW-13
Table C: Forgetting metrics in terms of mAP for the ODinW-13 dataset. The numbers above each task indicate the order in which they were processed. Bold indicates best results.
$11$ $10$ $2$ $13$ $12$ $1$ $7$ $6$ $3$ $4$ $9$ $8$ $5$
Method Avg Ae Aq Co Eg Mu Pa Pv Pi Po Ra Sh Th Ve
ZiRa -$\texttt{-}6.66$$\texttt{-}4.31$$\texttt{-}4.82$$\boldsymbol{\texttt{-}1.80}$ $0.00$$\texttt{-}2.09$$\texttt{-}8.71$$\texttt{-}5.80$$\texttt{-}7.76$$\texttt{-}12.63$$\texttt{-}19.19$$\boldsymbol{\texttt{-}1.82}$$\texttt{-}9.85$$\texttt{-}7.77$
DitHub -$\boldsymbol{\texttt{-}1.73}$$\boldsymbol{\texttt{-}1.56}$$\boldsymbol{\texttt{-}1.69}$$\texttt{-}2.34$ $0.00$$\boldsymbol{\texttt{-}1.61}$ $\boldsymbol{~{}~{}3.62}$$\boldsymbol{\texttt{-}1.91}$ $\boldsymbol{~{}~{}1.54}$$\boldsymbol{\texttt{-}6.94}$$\boldsymbol{\texttt{-}7.27}$$\texttt{-}2.11$ $\boldsymbol{~{}~{}0.10}$$\boldsymbol{\texttt{-}2.34}$
Tab. C presents the forgetting in terms of mAP points for the $13$ tasks of ODinW-13, with the average across all tasks reported as Avg. DitHub demonstrates strong performance on both average and individual tasks, consistently outperforming its competitor by a notable margin. Given the limited class overlap within ODinW-13, the primary component susceptible to catastrophic forgetting in this analysis is the shared matrix $B$.
Notably, for some tasks, this metric is even positive for our approach. We attribute this behavior to the initialization of the $B$ component. Specifically, the metric is higher in the first learned task (Pa)  [^4] The order in which tasks are presented to the model is random and dependent on the selected seed.. While this might seem counterintuitive, the Pa dataset comprises very few instances, which means the $B$ matrix cannot undergo robust initialization. Consequently, subsequent tasks introduce greater data diversity, leading to improved performance compared to the initial one limited by scarce training data.
B.2 Forgetting on ODinW-O
Table D: Forgetting metrics in terms of mAP for the ODinW-O dataset. Bold indicates best results.
$5$ $6$ $1$ $2$ $4$ $3$
Method Avg Ae Hw Pv Sd Th Ve
ZiRa -$\texttt{-}3.84$$\boldsymbol{\texttt{-}0.08}$ $0.00$$\texttt{-}9.67$$\texttt{-}10.66$$\texttt{-}0.98$$\texttt{-}1.62$
DitHub -$\boldsymbol{\texttt{-}2.68}$$\texttt{-}4.75$ $0.00$$\boldsymbol{\texttt{-}2.74}$$\boldsymbol{\texttt{-}9.51}$ $\boldsymbol{~{}~{}1.74}$$\boldsymbol{\texttt{-}0.79}$
Tab. D reports the forgetting metric results for the ODinW-O benchmark. Unlike the previous analysis, this assessment allows us to evaluate the forgetting of DitHub’s class-specific $A$ components for all classes within the dataset. Indeed, ODinW-O is specifically designed to test Incremental Learning capabilities, featuring complete class overlap across tasks. This overlap leads to catastrophic forgetting of the $A_{c}$ expert component for the same class $c$ across different domains. It’s worth noting that even in this setting, which emphasizes the $A$ component, the forgetting of $B$ remains a concern. Even in this challenging context, DitHub surpasses its competitor by an appreciable margin, demonstrating difficulty only with the Ae task and highlighting the non-trivial nature of merging and preserving knowledge from the aerial domain.
Appendix C Additional Benchmark Details